OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 3x+5y=3 x+2y=0

OpenStudy (anonymous):

solve the second equation for x

OpenStudy (anonymous):

x=-2y

OpenStudy (anonymous):

then 3x-5(-2y)=3

OpenStudy (anonymous):

isnt it 3(-2y)-5y=3?

OpenStudy (anonymous):

oh yeah it is

OpenStudy (anonymous):

-6y-5y=3 -11y=3 ---- --- -11 -11 y=3/-11

OpenStudy (anonymous):

But my choices are (6,1) (1,-3) (6,-3) (1,1)

OpenStudy (anonymous):

it is (6,-3)

OpenStudy (anonymous):

if you dont mind my asking, why?

OpenStudy (anonymous):

okay,; 3(-2y)+5y=3 -6y+5y=3 -1y=3 y=-3

OpenStudy (anonymous):

i see! Thank you